# Efficient Statistical Estimation for Sequential Adaptive Experiments with Implications for Adaptive Designs

### v1

@@ Abstract
Adaptive experimental designs have gained popularity in clinical trials and online experiments. Unlike traditional, fixed experimental designs, adaptive designs can dynamically adjust treatment randomization probabilities and other design features in response to data accumulated sequentially during the experiment. These adaptations are useful to achieve diverse objectives, including reducing uncertainty in the estimation of causal estimands or increasing participants’ chances of receiving better treatments during the experiment. At the end of the experiment, it is often desirable to answer causal questions from the observed data. However, the adaptive nature of such experiments and the resulting dependence among observations pose significant challenges to providing valid statistical inference and efficient estimation of causal estimands. Building upon the Targeted Maximum Likelihood Estimator (TMLE) framework tailored for adaptive designs (van der Laan 2008), we introduce a new adaptive-design-likelihood-based TMLE (ADL-TMLE) to estimate a variety of causal estimands from adaptive experiment data. We establish asymptotic normality and semiparametric efficiency of ADL-TMLE under relaxed positivity and design stabilization assumptions for adaptive experiments. Motivated by efficiency results, we further propose a novel adaptive design aimed at minimizing the variance of estimators based on data generated under that design. Using the average treatment effect as a representative example, simulation studies show that ADL-TMLE demonstrates superior variance-reduction performance across different types of adaptive experiments, and that the proposed adaptive design attains lower variance than the standard efficiency-oriented adaptive design. Finally, we generalize this estimation and design framework to broader settings with longitudinal structures.
@@ 1 Introduction
While adaptive experiments have advantage of improving estimation efficiency and/or bringing benefits to participants compared with those with fixed, non-adaptive designs, they also present challenges for statistical inference. At the end of an adaptive experiment, investigators often seek to estimate causal estimands based on the observed data, such as the average treatment effect or mean outcome under personalized treatment rules. However, the adaptivity of such experiments induces dependence among observations, violating the i.i.d. assumptions that standard statistical inference methods rely on. This has motivated substantial methodological work to provide statistical inference for data collected under adaptive designs. For example, Zhang et al. (2007) and Zhu (2015) derived asymptotic results for inference under CARA designs within parametric modeling frameworks. van der Laan (2008) introduced Targeted Maximum Likelihood Estimation (TMLE) framework that is tailored for adaptive design data and provides non-parametric statistical inference method without relying on parametric assumptions. Subsequent developments include the development of statistical inference for causal estimands such as the risk difference and log-relative risk (Chambaz & van der Laan 2014) as well as for estimating the mean outcome under optimal treatment rules (Murphy 2003, Luedtke & van der Laan 2016b) in CARA experiments (Chambaz et al. 2017). Another line of work employs variants of stabilized reweighting techniques (Luedtke & van der Laan 2016a) and develops asymptotically normal estimators for causal estimands in non‑contextual (Hadad et al. 2021) and contextual bandit settings (Bibaut et al. 2021, Zhan et al. 2021).
In this work, we propose ADL-TMLE, an adaptive-design-likelihood-based TMLE framework that estimates various causal estimands from adaptive experiment data by solving the score equation of the components that define the estimand. We establish both asymptotic normality and semiparametric efficiency of the new TMLE, and demonstrate relaxed positivity and stabilization conditions for adaptive experiment, along with its superior finite-sample efficiency compared with the previous adaptive-design TMLE. Motivated by these efficiency results, we further develop a novel adaptive design framework aimed at minimizing variance of efficient estimators built upon the resulting adaptive experiment data. Finally, we generalize our estimation and design framework to broader settings with longitudinal structures.
@@
The remainder of the paper is organized as follows. Section 2 introduces the statistical setup. Section 3 provides the estimation procedure of the proposed TMLE using ATE as illustrating example. Sections 4 and 5 establish the asymptotic normality and efficiency properties. Section 6 presents the proposed adaptive design approach. Section 7 reports simulation studies evaluating the performance of the proposed estimation and design methods. Section 8 discusses applications and generalizations of our framework, followed by conclusions in Section 9.
@@ 2 Data, Model and Statistical Setup
We define the full data as $X=(W,Y(1),Y(0))\sim P^{F}\in\mathcal{M}^{F}$, where $W\in\mathcal{W}$ are baseline covariates, $Y(a)\in\mathcal{Y}$ is the counterfactual outcome under treatment level $a\in\mathcal{A}$. Without loss of generality, we consider binary treatment space $\mathcal{A}=\{0,1\}$ and set $\mathcal{Y}=[0,1]$. Here $P^{F}$ is the true distribution that generates full data $X$, living in a general model $\mathcal{M}^{F}$ without model prespecification. We use $E_{P^{F}}$ to denote the expectation under $P^{F}$. While our framework is generally applicable for estimating a wide class of causal estimands (see Section 8), here we use average treatment effect (ATE) $\Psi(P^{F}):=E_{P^{F}}[Y(1)-Y(0)]$ as the main example for illustration.
We consider an adaptive experiment that enrolls $n$ units sequentially, where $i$ is an index of each unit reflecting their enrollment order and $O_{i}:=(W_{i},A_{i},Y_{i})$ are the data to be observed for this unit. Upon the enrollment, the experimenter first observes baseline covariates $W_{i}$, then assigns treatment $A_{i}$ following a treatment randomization function $g_{i}:\mathcal{W}\times\mathcal{A}\to[0,1]$ within a function class $\mathcal{G}$. Then, the unit’s outcome $Y_{i}$ is observed. The adaptivity of the experiment is reflected in that treatment $A_{i}$ is determined by a function $g_{i}$ that is itself generated based on the observed data of previously enrolled units before $i$, which we denote by $\bar{O}(i-1):=(O_{1},\cdots,O_{i-1})$. We use $\mathbf{g}_{1:n}:=(g_{1},\cdots,g_{i},\cdots,g_{n})$ to refer to an adaptive design that includes a sequence of treatment randomization functions $g_{i}\in\mathcal{G}$, which is known and controlled by the experimenter.
Define $\mathcal{Q}:=\mathcal{Q}_{W}\otimes\mathcal{Q}_{Y}$, where $\mathcal{Q}_{W}$ is the set of possible distributions of $W$ and $\mathcal{Q}_{Y}$ is the set of possible conditional distributions of $Y$ given $A$ and $W$. For any $Q=(Q_{W},Q_{Y})\in\mathcal{Q}$, we denote $\bar{Q}(A,W)$ as a shorthand of $E[Y|A,W]$, the conditional mean of $Y$ given $A$ and $W$. For every $i\in[n]$, we assume that $W_{i}\sim Q_{W}$ and $Y_{i}|A_{i},W_{i}\sim Q_{Y}$. We use $P_{Q}^{n}\in\mathcal{M}^{n}$ to denote the distribution that generates the adaptive experiment data $\bar{O}(n)$ in the adaptive design model $\mathcal{M}^{n}$. The model specifies the likelihood of $\bar{o}(n)=(o_{1},\cdots,o_{i},\cdots,o_{n})$ as
@@
One can identify the causal estimand $\Psi(P^{F})$ defined on full data $P^{F}$ from a target estimand $\psi(Q)$ with the following assumptions:
@@
Theorem 1 (Identification).
With Assumptions 1 and 2, one can identify $\Psi(P^{F})$ from the observed data $\bar{O}(n)$ by
$\Psi(P^{F})$ $=$ $E_{Q_{W}}E[\bar{Q}(1,W)-\bar{Q}(0,W)]$
$:=$ $\psi(Q).$
@@ 3 Targeted Maximum Likelihood Estimation
At the end of an adaptive experiment, one observes a single series of dependent data $\bar{O}(n)=(O_{1},\cdots,O_{n})$. Let $\bar{Q}_{n}$ denote an initial estimate of the true conditional mean function $\bar{Q}_{0}$ based on either a user-specified model or an ensemble learner such as Super Learner (van der Laan et al. 2007). For any adaptive design $\mathbf{g}_{1:n}$, we define its average design $\bar{g}_{n}:=\frac{1}{n}\sum_{i=1}^{n}g_{i}$. Our proposed ADL-TMLE is
@@
where $Q_{n}^{*}=(Q_{n,W},Q_{n,Y}^{*})$ where $Q_{n,W}$ is the empirical distribution of baseline covariates, and $Q_{n,Y}^{*}$ is an updated outcome distribution associated with a TMLE update $\bar{Q}_{n}^{*}$ of the initial estimate $\bar{Q}_{n}$. Specifically, $\bar{Q}_{n}^{*}$ is obtained by solving
@@
via the regression: $\mathop{\mathrm{logit}}\bar{Q}_{n,\epsilon}\left(A,W\right)=\mathop{\mathrm{logit}}\bar{Q}_{n}\left(A,W\right)+\epsilon H_{n}(A,W)$, where $H_{n}$ is defined by $H_{n}(A,W):=\frac{2A-1}{\bar{g}_{n}(A|W)}$ based on the average design $\bar{g}_{n}$. The TMLE update is $\bar{Q}_{n}^{*}=\bar{Q}_{n,\epsilon_{n}}$, where $\epsilon_{n}$ is the solution of that regression.
@@
We note that our proposed ADL-TMLE is driven by the likelihood structure of adaptive experiment and solves the score equations for the $Q$ components that define the target parameter $\psi(Q)$. For any $Q=(Q_{W},Q_{Y})\in\mathcal{Q}$ with the associated conditional mean function $\bar{Q}\in\bar{\mathcal{Q}}$ and a treatment randomization function $g\in\mathcal{G}$, define $D(Q,g)$ as the efficient influence curve (EIC) of the target estimand in the i.i.d. experiment: $W\sim Q_{W},A|W\sim g$, and $Y|A,W\sim Q_{Y}$. For the ATE $\psi(Q)$, the EIC is $D(Q,g)(O)=D_{1}(\bar{Q},g)(O)+D_{2}(\bar{Q})(O)$, where $D_{1}(\bar{Q},g)(O):=\frac{2A-1}{\bar{g}(A|W)}\left(Y-\bar{Q}(A,W)\right)$ and $D_{2}(Q)(O):=\bar{Q}(1,W)-\bar{Q}(0,W)-\psi(Q)$. The core procedure of ADL-TMLE is to solve $\frac{1}{n}\sum_{i=1}^{n}D(Q,\bar{g}_{n})=0$, which behaves as solving the score equation in the i.i.d. experiment with $A\sim\bar{g}_{n}$, even when the data is collected from the adaptive experiment. This framework is general for estimating a wider class of causal estimands as shown in Section 8.
For comparison, we also introduce the original TMLE tailored for adaptive design data proposed by van der Laan (2008), denoted by $\psi(Q_{n}^{\dagger})$ and referred to as “AD-TMLE” (adaptive-design TMLE). In AD-TMLE, $Q_{n}^{\dagger}=(Q_{n,W},Q_{n,Y}^{\dagger})$ and $\bar{Q}_{n}^{\dagger}$ is the TMLE update of the estimated conditional mean function $\bar{Q}_{n}$, which solves $\frac{1}{n}\sum_{i=1}^{n}\frac{2A_{i}-1}{g_{i}(A_{i}|W_{i})}[Y_{i}-\bar{Q}_{n}^{\dagger}(A_{i},W_{i})]=0$ using every $i$-specific treatment randomization function $g_{i}$ from the adaptive design $\mathbf{g}_{1:n}$. Its variance estimator is given by $\tilde{\sigma}_{n}^{2}/n$, where $\tilde{\sigma}_{n}^{2}:=\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{2A_{i}-1}{g_{i}(A_{i}|W_{i})}\left[Y_{i}-\bar{Q}_{n}^{\dagger}(A_{i},W_{i})\right]+\bar{Q}_{n}^{\dagger}(1,W_{i})-\bar{Q}_{n}^{\dagger}(0,W_{i})-\psi(Q_{n}^{\dagger})\right\}^{2}$. While the next sections focus on the properties of TMLEs, these analyses also apply to the corresponding One-Step/AIPW estimators $\hat{\psi}_{n}(\bar{Q}_{n}):=\frac{1}{n}\sum_{i=1}^{n}\frac{2A_{i}-1}{\bar{g}_{n}(A_{i}|W_{i})}\left[Y_{i}-\bar{Q}_{n}(A_{i},W_{i})\right]+\frac{1}{n}\sum_{i=1}^{n}\left[\bar{Q}_{n}(1,W_{i})-\bar{Q}_{n}(0,W_{i})\right]$ and $\hat{\psi}_{n}^{\dagger}(\bar{Q}_{n}):=\frac{1}{n}\sum_{i=1}^{n}\frac{2A_{i}-1}{g_{i}(A_{i}|W_{i})}\left[Y_{i}-\bar{Q}_{n}(A_{i},W_{i})\right]+\frac{1}{n}\sum_{i=1}^{n}\left[\bar{Q}_{n}(1,W_{i})-\bar{Q}_{n}(0,W_{i})\right].$

### v2

@@ Abstract
Adaptive experimental designs have gained popularity in clinical trials and online experiments. Unlike traditional, fixed experimental designs, adaptive designs can dynamically adjust treatment randomization probabilities and other design features in response to data accumulated sequentially during the experiment. These adaptations are useful to achieve diverse objectives, including reducing uncertainty in the estimation of causal estimands or increasing participants’ chances of receiving better treatments during the experiment. At the end of the experiment, it is often desirable to answer causal questions from the observed data. However, the adaptive nature of such experiments and the resulting dependence among observations pose significant challenges to providing valid statistical inference and efficient estimation of causal estimands. Building upon the Targeted Maximum Likelihood Estimation (TMLE) framework tailored for adaptive designs (van der Laan 2008), we introduce a new Adaptive-Design-Likelihood-based TMLE (ADL-TMLE) to estimate a wide class of causal estimands from adaptive experiment data, including the average treatment effect as our primary example. We establish asymptotic normality and semiparametric efficiency of ADL-TMLE under relaxed positivity and design stabilization assumptions for adaptive experiments. Motivated by these results, we further propose a novel adaptive design aimed at minimizing the variance of the estimator based on data generated under that design. Simulations show that ADL-TMLE demonstrates superior variance-reduction performance across different types of adaptive experiments, and that the proposed adaptive design attains lower variance than the standard efficiency-oriented adaptive design. Finally, we generalize our framework to broader settings, including those with longitudinal structures.
@@ 1 Introduction
While adaptive experiments have advantage of improving estimation efficiency and/or bringing benefits to participants compared with those with fixed, non-adaptive designs, they also present challenges for statistical inference. At the end of an adaptive experiment, investigators often seek to estimate causal estimands based on the observed data, such as the average treatment effect or mean outcome under personalized treatment rules. However, the adaptivity of such experiments induces dependence among observations, violating the i.i.d. assumptions that standard statistical inference methods rely on. This has motivated substantial methodological work to provide statistical inference for data collected under adaptive designs. For example, Zhang et al. (2007) and Zhu (2015) derived asymptotic results for inference under CARA designs within parametric modeling frameworks. van der Laan (2008) introduced Targeted Maximum Likelihood Estimation (TMLE) framework which is tailored for adaptive design data and provides non-parametric statistical inference method without relying on parametric assumptions. Subsequent developments include the development of statistical inference for causal estimands such as the risk difference and log-relative risk (Chambaz & van der Laan 2014) as well as for estimating the mean outcome under optimal treatment rules (Murphy 2003, Luedtke & van der Laan 2016b) in CARA experiments (Chambaz et al. 2017). Another line of work employs variants of stabilized reweighting techniques (Luedtke & van der Laan 2016a) and develops asymptotically normal estimators for causal estimands in non‑contextual (Hadad et al. 2021) and contextual bandit settings (Bibaut et al. 2021, Zhan et al. 2021).
In contrast to these methods, we develop a unified framework that enables semiparametrically efficient estimation and inference for adaptive experiments and provides a new adaptive design that targets estimation efficiency.
First, we introduce a novel Adaptive-Design-Likelihood-based TMLE (ADL-TMLE) framework that enables statistical estimation and inference for a wide class of causal estimands from adaptive experiments, including the average treatment effect as our primary example. This estimator is driven by the likelihood structure of adaptive experiment and solves the score equation for components that exclusively define the causal estimand of interest. We show that ADL-TMLE is doubly robust, asymptotically normal and semiparametrically efficient.
A key insight from these results is that the minimal asymptotic variance of regular estimators under the adaptive experiment is inherently linked to that under a reference i.i.d. experiment in which treatments are assigned by the average of treatment randomization functions across the adaptive experiment (which we refer to as the “average design”). This highlights the advantage of ADL-TMLE in enabling statistical inference under relaxed assumptions for adaptive experiments. Specifically, it suffices to require positivity only for the average design, rather than for every individual treatment assignment probability. This enables greater flexibility in adaptive experiments, where local positivity violations are allowed during some segments of the experiment. The design stabilization conditions for establishing asymptotic normality and efficiency are also relaxed, as they are required only for the average design, which also leads to improved finite-sample performance of ADL-TMLE. Moreover, in scenarios where individual-level treatment randomization mechanisms are inaccessible due to operational or privacy constraints, our estimation framework remains feasible as long as the average design is accessible. These theoretical and practical advantages are further supported by simulation studies, which demonstrate the superior variance-reduction performance of ADL-TMLE across different adaptive experiments with varied objectives.
Second, while our framework provides a semiparametrically efficient estimator for a given adaptive experiment, even greater efficiency can be attained by guiding the real-time adaptive design toward an oracle design that is itself optimized for estimation efficiency. Motivated by ADL-TMLE framework, we propose a novel adaptive design framework that dynamically steers the average design toward the evolving estimate of the oracle design as data accumulate throughout the experiment. This adaptive procedure leads to a lower variance than the standard efficiency-oriented adaptive design as shown by simulations.
Finally, our framework is directly applicable to settings with independent data structures, such as cases where certain factors induce heterogeneity in the treatment assignment mechanism but do not affect the primary outcome conditional on treatment and other baseline covariates. Furthermore, we generalize our framework to broader settings, including those with longitudinal data structures.
@@
The remainder of the paper is organized as follows. Section 2 introduces the statistical setup. Section 3 provides the estimation procedure of ADL-TMLE. Sections 4 and 5 establish asymptotic normality and semiparametric efficiency. Section 6 describes the proposed adaptive design approach. Section 7 reports simulation studies evaluating the performance of the proposed estimation and design methods. Section 8 discusses applications and generalizations of our framework. Finally, Section 9 concludes.
@@ 2 Data, Model and Statistical Setup
We define the full data as $X=(W,Y(1),Y(0))\sim P^{F}\in\mathcal{M}^{F}$, where $W\in\mathcal{W}$ are baseline covariates, $Y(a)\in\mathcal{Y}$ is the counterfactual outcome under treatment level $a\in\mathcal{A}$. Without loss of generality, we consider binary treatment space $\mathcal{A}=\{0,1\}$ and set $\mathcal{Y}=[0,1]$. Here $P^{F}$ is the true distribution that generates full data $X$, living in a general model $\mathcal{M}^{F}$ without model prespecification. We use $E_{P^{F}}$ to denote the expectation under $P^{F}$. While our framework is generally applicable for estimating a wide class of causal estimands (see Section 8), here we use ATE $\Psi(P^{F}):=E_{P^{F}}[Y(1)-Y(0)]$ as the main example for illustration.
Consider an adaptive experiment that enrolls $n$ units sequentially, where $i$ is an index of each unit reflecting their enrollment order and $O_{i}:=(W_{i},A_{i},Y_{i})$ are the data to be observed for this unit. Upon the enrollment, the experimenter first observes baseline covariates $W_{i}$, then assigns treatment $A_{i}$ following a treatment randomization function $g_{i}:\mathcal{W}\times\mathcal{A}\to[0,1]$ within a function class $\mathcal{G}$. Then, the unit’s outcome $Y_{i}$ is observed. The adaptivity of the experiment is reflected in that treatment $A_{i}$ is determined by a function $g_{i}$ that is itself generated based on the observed data of previously enrolled units before $i$, which we denote by $\bar{O}(i-1):=(O_{1},\cdots,O_{i-1})$. We use $\mathbf{g}_{1:n}:=(g_{1},\cdots,g_{i},\cdots,g_{n})$ to refer to an adaptive design that includes a collection of treatment randomization functions $g_{i}\in\mathcal{G}$ from $i=1,\cdots,n$, which are known and controlled by the experimenter.
Define $\mathcal{Q}:=\mathcal{Q}_{W}\otimes\mathcal{Q}_{Y}$, where $\mathcal{Q}_{W}$ is the set of possible distributions of $W$ and $\mathcal{Q}_{Y}$ is the set of possible conditional distributions of $Y$ given $A$ and $W$. For any $Q=(Q_{W},Q_{Y})\in\mathcal{Q}$, we denote $\bar{Q}(A,W)$ as a shorthand of $E[Y|A,W]$, the conditional mean of $Y$ given $A$ and $W$. For every $i\in[n]$, we assume that $W_{i}\sim Q_{W}$ and $Y_{i}\sim Q_{Y}(\cdot|A_{i},W_{i})$. We use $P_{Q}^{n}\in\mathcal{M}^{n}$ to denote the distribution that generates the adaptive experiment data $\bar{O}(n)$ in the adaptive design model $\mathcal{M}^{n}$. The model specifies the likelihood of $\bar{o}(n)=(o_{1},\cdots,o_{i},\cdots,o_{n})$ as
@@
One can identify the causal estimand $\Psi(P^{F})$ with the following assumptions:
@@
Theorem 1 (Identification of ATE).
With Assumptions 1 and 2, $\Psi(P^{F})$ is identified by
$\Psi(P^{F})$ $=$ $E[\bar{Q}(1,W)-\bar{Q}(0,W)].$
We denote this target estimand by $\psi(Q):=E[\bar{Q}(1,W)-\bar{Q}(0,W)]$.
@@ 3 Targeted Maximum Likelihood Estimation
At the end of an adaptive experiment, one observes a single series of dependent data $\bar{O}(n)=(O_{1},\cdots,O_{n})$. Let $\bar{Q}_{n}$ denote an initial estimate of the true conditional mean function $\bar{Q}_{0}$, estimated by a user-specified model or an ensemble learner such as Super Learner (van der Laan et al. 2007). For any adaptive design $\mathbf{g}_{1:n}$, we define its average design $\bar{g}_{n}:=\frac{1}{n}\sum_{i=1}^{n}g_{i}$.
Our proposed ADL-TMLE is
@@
where $Q_{n}^{*}=(Q_{n,W},Q_{n,Y}^{*})$ with $Q_{n,W}$ denotes the empirical distribution of baseline covariates, and $Q_{n,Y}^{*}$ an updated outcome distribution associated with a TMLE update $\bar{Q}_{n}^{*}$ of the initial estimate $\bar{Q}_{n}$. Specifically, $\bar{Q}_{n}^{*}$ is obtained by solving
@@
via the following regression:
$\mathop{\mathrm{logit}}\bar{Q}_{n,\epsilon}\left(A,W\right)=\mathop{\mathrm{logit}}\bar{Q}_{n}\left(A,W\right)+\epsilon H_{n}(A,W),$
where $H_{n}$ is defined by $H_{n}(A,W):=\frac{2A-1}{\bar{g}_{n}(A|W)}$ based on the average design $\bar{g}_{n}$. The TMLE update is $\bar{Q}_{n}^{*}=\bar{Q}_{n,\epsilon_{n}}$, where $\epsilon_{n}$ is the solution of that regression.
@@
We note that our proposed ADL-TMLE is driven by the likelihood structure of adaptive experiment and solves the score equation for the $Q$-components that exclusively define the target parameter $\psi(Q)$. For any $Q=(Q_{W},Q_{Y})\in\mathcal{Q}$ and treatment randomization function $g\in\mathcal{G}$, we denote $D(Q,g)$ as the efficient influence curve (EIC) of the target estimand in an i.i.d. experiment where $W\sim Q_{W},A\sim g(\cdot|W)$, and $Y\sim Q_{Y}(\cdot|A,W)$. For ATE, the EIC is $D(Q,g)(O)=D_{1}(Q,g)(O)+D_{2}(Q)(O),$ where $D_{1}(Q,g)(O):=\frac{2A-1}{g(A|W)}\left(Y-\bar{Q}(A,W)\right)$ and $D_{2}(Q)(O):=\bar{Q}(1,W)-\bar{Q}(0,W)-\psi(Q)$. The core procedure of ADL-TMLE is to solve $\frac{1}{n}\sum_{i=1}^{n}D(Q,\bar{g}_{n})(O_{i})=0$, which is equivalent to solving the score equation as if in the i.i.d. experiment where the treatment distribution follows $A\sim\bar{g}_{n}(\cdot|W)$, even though the data is collected from the adaptive experiment.
For comparison, we also introduce the original TMLE tailored for adaptive design data proposed by van der Laan (2008), which we denote by $\psi(Q_{n}^{\dagger})$ and refer to as “AD-TMLE” (adaptive-design TMLE). In AD-TMLE, $Q_{n}^{\dagger}=(Q_{n,W},Q_{n,Y}^{\dagger})$, and $\bar{Q}_{n}^{\dagger}$ is the TMLE update of the estimated conditional mean function $\bar{Q}_{n}$, which solves $\frac{1}{n}\sum_{i=1}^{n}\frac{2A_{i}-1}{g_{i}(A_{i}|W_{i})}[Y_{i}-\bar{Q}_{n}^{\dagger}(A_{i},W_{i})]=0$ using every $i$-specific treatment randomization function $g_{i}$ from the adaptive design $\mathbf{g}_{1:n}$. Its variance estimator is given by $\tilde{\sigma}_{n}^{2}/n$, where $\tilde{\sigma}_{n}^{2}:=\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{2A_{i}-1}{g_{i}(A_{i}|W_{i})}\left[Y_{i}-\bar{Q}_{n}^{\dagger}(A_{i},W_{i})\right]+\bar{Q}_{n}^{\dagger}(1,W_{i})-\bar{Q}_{n}^{\dagger}(0,W_{i})-\psi(Q_{n}^{\dagger})\right\}^{2}$. While next sections focus on properties of TMLEs, they also apply to the corresponding One-Step/AIPW estimators $\hat{\psi}_{n}(\bar{Q}_{n}):=\frac{1}{n}\sum_{i=1}^{n}\frac{2A_{i}-1}{\bar{g}_{n}(A_{i}|W_{i})}\left[Y_{i}-\bar{Q}_{n}(A_{i},W_{i})\right]+\frac{1}{n}\sum_{i=1}^{n}\left[\bar{Q}_{n}(1,W_{i})-\bar{Q}_{n}(0,W_{i})\right],$ and $\hat{\psi}_{n}^{\dagger}(\bar{Q}_{n}):=\frac{1}{n}\sum_{i=1}^{n}\frac{2A_{i}-1}{g_{i}(A_{i}|W_{i})}\left[Y_{i}-\bar{Q}_{n}(A_{i},W_{i})\right]+\frac{1}{n}\sum_{i=1}^{n}\left[\bar{Q}_{n}(1,W_{i})-\bar{Q}_{n}(0,W_{i})\right]$.